OpenStudy (anonymous):

Can someone please help me solve this function? Given f(x)=5x-8 and g(x)=x^2+4 Find (fog)(1) and (gof)(x)

jimthompson5910 (jim_thompson5910):

(fog)(x) = f(g(x)) (fog)(x) = f(x^2+4) (fog)(1) = f(1^2+4) (fog)(1) = f(1+4) (fog)(1) = f(5) (fog)(1) = 5(5)-8 (fog)(1) = 25-8 (fog)(1) = 17